Problem - Determine the total number of energy states in silicon between E, and E + kT at T = 300 K.

1
Expert's answer
2022-02-01T15:33:56-0500

"n=2(\frac{2\pi m^* kT}{h^2})^\frac 32=2.81\cdot 10^{25}~m^{-3}."